Hosts A and B are each connected to a switch S via 1-Gbps links as below. The propagation delay on each link is 20 μsHoOi5bySvnhWFUZIIzqsGlxsyktyouB0JdXK333Q. S is a store-and-forward device; it begins retransmitting a received packet 35μsHoOi5bySvnhWFUZIIzqsGlxsyktyouB0JdXK333Q after it has finished receiving it. Calculate the total time required to transmit 10,000 bits from A to B in the following two cases:

  • When all the 10,000 bits are transmitted in a single packet?
  • When the 10,000 bits are transmitted in two 5,000 bits packets sent one right after the other.
Answer #1

Link Bandwidth = 1 Gbps

Propagation delay = 20 micro seconds

Answer 1

Single packet will take time = number of links* (Propagation time + Transmission time) + switch delay

= 20 \mu s + (10000/ 109) * 106 \mu s + 35 \mu s

= 2*(20 \mu s + 10 \mu s) + 35 \mu s

= 95 \mu s

Answer 2

Both packets will be transmitted after each other.

First bit of first packet will reach to switch in 20 \mu s

First complete packet of 5000 bits will be received at switch S in

20 \mu s + (5000/109) * 106 \mu s = 25 \mu s

Switch S will start re-transmission after 35 \mu s and Host B will start receiving packet B (first bit) after 20 \mu s

25 \mu s + 35 \mu s + 20 \mu s = 80 \mu s

Meanwhile packet 2 has started transmission.

Similarly second packet of 5000 bits will be received at switch S in

= time when first packet is transmitted fully + transmission time for second packet = 25 \mu s + 5 \mu s = 30 \mu s

Packet 2 will wait at switch S till first packet is not transmitted fully. S will finish first packet re-transmission at time = 60 \mu s + 5 \mu s = 65 \mu s.

S will start sending second packet at 65 \mu s. Host B will receive full first packet after 80 \mu s + 5 \mu s = 85 \mu s. Host B will receive second packet after 85 \mu s + 5 \mu s = 90 \mu s

So two packets take lesser time than single packet.